# Hierarchical Refinement of Universal Multimodal Attacks on Vision-Language Models

### v1

@@ III-A Preliminaries
The proposed HRA learns universal adversarial perturbations (UAPs) for both the image and text modalities in the black-box setting, where the target VLP models, datasets, and downstream tasks are unknown and inaccessible during training. To tackle this challenge, we introduce a transfer learning-based attack method that leverages a surrogate model and dataset to generate UAPs capable of transferring across diverse models, datasets, and tasks.
Our approach is generic and does not impose specific assumptions about the surrogate model used during training. For example, some VLP models employ a multimodal encoder [16, 39], while others do not [29]. To ensure broad generality, we uniformly utilize the unimodal encoders available in each VLP model, i.e., the image encoder and the text encoder, denoted as and , respectively. Suppose a multi-modal dataset is available for training, where is an image-text pair and is the number of pairs. The objective is to learn UAPs that can mislead other VLP models to incorrectly associate images and texts in a target dataset at the reference time. For the image modality, we learn UAPs constrained by an norm to ensure imperceptibility, i.e., . Here, represents the perturbation magnitude. For the text modality, due to token discreteness, we implement a universal attack by consistently substituting a word in the text with a predefined replacement word. To maintain imperceptibility, the number of such substitutions is limited by a substitution budget .
III-A1 Vanilla Learning
VLP models commonly bridge the gap between heterogeneous data, i.e., images and texts, by creating a shared representation space to associate images and texts with similar semantics. This makes it natural for an attacker to disturb the alignment between paired images and texts. Current methods usually only concentrate on the image modality. Specifically, given a surrogate model, UAPs for the image modality are learned by enlarging the feature distance between matched images and texts:
@@ III-A1 Vanilla Learning
Iterative methods such as PGD [14, 25] are usually used to solve the above objective. This scheme can effectively work in white-box scenarios, where data and models to attack during training are the same as those during testing. However, when applying learned UAPs to attack other different models, data, and tasks, i.e., black-box settings, UAPs would lose efficacy. We own this failure to several reasons. First, this vanilla UAP learning only enlarges the feature distance between surrogate image-text pairs on the source model, without considering the generalization of UPAs to other models, tasks and datasets. Different VLP models have distinct architectures, training objectives and datasets across both pre-training and fine-tuning stages. These variations cause different models to rely on different features and patterns to associate modalities and make decisions. For example, image-text retrieval retrieves relevant data for a given query by ranking their similarity in the feature space, while Image captioning generates descriptive text for an image by recognizing objects, understanding scenes, and capturing relationships within the image. Attacking retrieval models requires disrupting ranking consistency, whereas misleading captioning models demands precise manipulation of semantic content. Second, vanilla UAP learning fails to account for the inherent heterogeneity between cross-modal data, limiting its ability to disrupt cross-modal alignment effectively. Since images and texts differ in format and provide diverse information, complex cross-modal interactions exist between them. As a result, simply increasing the feature distance between images and texts does not ensure maximal cross-modal misalignment. Lastly, UAPs are optimized over an entire dataset, requiring them to be robust against diverse sample patterns. To achieve this, the learning process often exploits shortcuts to the source model, generating perturbations that are highly specific to it. As a result, it leads to severe overfitting, reducing generalizability to unseen models and tasks. To enable effective and transferable attacks, it is crucial to improve cross-modal alignment disruption while reducing dependence on specific models, tasks, and data patterns.
III-B Overview
To handle sample-, model- and task-independent challenges, we propose universal multimodal adversarial attacks with hierarchical refinement (HRA), which optimizes UAPs from two perspectives: cross-modal alignment and the optimization path. For the UAP learning of the image modality, HRA disentangles adversarial examples to separately deal with original data and UAPs. For original data, a novel ScMix data augmentation and benign adversarial perturbations are leveraged to increase data diversity, facilitating the exploration of cross-modal alignment and reducing reliance on specific data features and patterns. Additionally, both global and local regions of UAPs are reinforced to reduce dependence on specific global features and prevent the formation of shortcuts, thereby mitigating overfitting. Furthermore, HRA directly regularizes the optimization path, as overfitting occurs when the optimization drops into local optima. By incorporating both past and future gradients from vanilla learning, it stabilizes update directions, preventing convergence to local optima. For the text modality, the proposed method ranks word importance by masking individual tokens and measuring the divergence between the masked and original text representations, considering both intra-sentence and inter-sentence contexts. An illustration of the proposed HRA can be found in Fig. 1.
III-C Disentangled Adversarial Images
| Test Dataset | Flickr30K | ||||||
| Task | Image-to-Text | Text-to-Image | |||||
| Target Model | Method | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 |
| HRA (image attack only) | 88.96 | 76.43 | 70.43 | 93.49 | 88.27 | 85.32 | |
| HRA (image attack only) | 93.13 | 88.16 | 83.84 | 96.13 | 93.83 | 92.37 | |
| CLIP | HRA (image attack only) | 88.47 | 78.61 | 73.07 | 92.69 | 87.5 | 84.89 |
| Test Dataset | MSCOCO | ||||||
| HRA (image attack only) | 29.21 | 14.32 | 9.57 | 27.79 | 14.35 | 10.28 | |
| HRA (image attack only) | 24.65 | 12.20 | 8.60 | 25.58 | 12.92 | 9.22 | |
| ALBEF | HRA (image attack only) | 32.43 | 15.76 | 11.28 | 29.82 | 16.25 | 11.83 |
III-C1 ScMix Augmentation.
Data augmentation is a well-established technique for improving adversarial transferability by increasing data diversity [24]. In multi-modal attack scenarios, employing data augmentation can help further exploit cross-modal interactions.
Specifically, an effective data transformation for augmentation should meet two key criteria: diversity and semantic preservation. This means the augmentation should introduce visual differences between the original and augmented data while maintaining semantic integrity. If the transformation significantly alters semantics, the model may fail to recognize the data correctly, ultimately reducing the effectiveness of the generated image UAPs. Under this regime, we propose a novel semantic-preserving ScMix method. ScMix incorporates self-mix and cross-mix operations. In the self-mix process, new data is constructed by mixing different transformed versions of the same original image. Specifically, two subregions of an image are randomly cropped and resized to match the original image size. Let and denote two rescaled subregions. They are then combined to create a new instance . Finally, the mixed instance is interpolated with another randomly selected image from the mini-batch to generate the final augmented sample . Formally, the whole process can be formulated as follows:
where ensures that the cross-mix does not significantly alter the original semantics when injecting another image. represents the prediction for the ScMixed data. denotes a Beta distribution with .
With ScMix, we can augment the original data to create diverse multi-modal data pairs. Furthermore, in datasets, e.g., Flickr30K [28] and MSCOCO [19], there often exists diverse captions that describe an image for different perspectives or language styles. We capture the most relevant captions for each image to further enhance multi-modal data diversity. Specifically, let the caption set for image be . The augmented data pairs are , where are generated by applying ScMix to . As a result, an enriched training set is obtained, which helps to better exploit cross-modal interactions and prevent overfitting.
Since UAPs are optimized across the entire dataset, applying data augmentation to increase data diversity for optimizing the full image UAPs may lead to overfitting. Table I presents some black-box attack results, where we observe that HRA that leverages the ScMix to optimize the entire UAP region, performs worse than HRA, which does not incorporate ScMix. To handle this, we propose leveraging ScMix to optimize local regions of image UAPs.
III-C2 Local Utility Enhancement.
Enhancing the local utility of image UAPs alongside optimizing the entire perturbation can prevent UAPs from relying on specific features and creating shortcuts to source models. Specifically, during the UAP learning process, we randomly crop subregions of UAPs and resize them to match the original image size. For brevity, we denote this transformation process as . Similar to optimizing the entire UAPs, the local regions of the UAPs are trained by enlarging the embedding distance between perturbed images and their original pairs:
where is the caption number.
The proposed local utility enhancement is effective for three key reasons. First, strengthening local regions naturally improves the overall effectiveness of image UAPs by making subregions more disruptive to various image patterns. This also facilitates disrupting cross-modal alignment across different samples, models and tasks, where diverse features are considered. Lastly, optimizing local regions acts as a form of model augmentation by introducing an extra transformation to image UAPs before feeding them into source models. This regularization effect helps mitigate overfitting, ensuring the perturbations remain effective across different scenarios.
III-C3 Overall objective.
The overall objective for optimizing UAPs for images is defined as:
where data in and would also be augmented with different captions for each image.
III-D Future-Aware Momentum
HRA further refines image UAPs by directly regularizing the optimization path. Gradient-based methods optimize UAPs by calculating the gradient direction of the loss Eq. (4). However, in multi-round learning procedures, the update process would gradually converge to local optimal, causing overfitting [8, 35, 33, 18]. The momentum method is an effective approach for stabilising update directions by leveraging previous gradients and has been widely used in various robust learning methods [23, 13]. Classical momentum [8] typically relies on accumulated past gradients to regularize the optimization trajectory. However, since UAPs are optimized across the entire dataset, gradients based on historical perturbations may diverge from the current optimization direction. Moreover, such momentum only consider a limited optimization scope, i.e., the current and past, resulting in a narrow exploration space. To address this, we propose Hierarchical Momentum that considers both past gradient and future gradients to stabilize the optimization process and broaden the search scope, thereby further alleviating overfitting.
@@ III-D Future-Aware Momentum
The previous gradient direction is less prone to overfitting, while the gradient direction from vanilla learning is more susceptible to overfitting. Considering this, we regularize the current gradient using both the previous and future gradients as follows:
@@ III-E Text Attack
III-F Algorithm
The detailed algorithm is summarized in Algorithm 1. For the image attack, in each epoch, mini-batch data are randomly sampled from the dataset. Next, ScMix is leveraged to create diverse data pairs. Subsequently, the current gradient, previous gradient, and future gradient are computed to update the UAPs. The optimization process repeats until the maximum number of iterations is reached. For the text attack, the method computes the importance of each word and aggregates the scores across the entire dataset, from which the most influential words are selected to serve as universal perturbations.
@@ IV Experiments
| Test Dataset | MSCOCO | ||||||||||||||
| AdvCLIP | 35.78 | 44.71 | 25.07 | 11.38 | 82.18 | 81.77 | 31.63 | 38.89 | 38.23 | 38.72 | 13.56 | 16.04 | 29.15 | 30.24 | |
| SGA | 96.49 | 95.62 | 56.03 | 59.45 | 96.49 | 95.62 | 41.55 | 46.68 | 36.44 | 41.41 | 23.25 | 23.45 | 24.74 | 24.15 | |
| ETU | 78.55 | 79.37 | 72.01 | 73.59 | 93.55 | 94.25 | 45.14 | 50.51 | 43.49 | 47.35 | 32.43 | 29.82 | 34.02 | 30.28 | |
| FD-UAP | 64.61 | 66.89 | 54.53 | 56.51 | 88.55 | 86.40 | 42.88 | 48.34 | 38.99 | 40.17 | 24.57 | 24.66 | 25.90 | 24.84 | |
| C-PGC | 87.29 | 89.36 | 82.71 | 86.80 | 95.76 | 95.85 | 74.40 | 80.89 | 69.13 | 77.25 | 54.68 | 55.39 | 72.41 | 76.33 | |
| HRAran | 87.37 | 91.84 | 86.11 | 91.74 | 96.38 | 97.01 | 65.01 | 80.05 | 78.02 | 88.06 | 53.49 | 68.76 | 58.92 | 72.54 | |
| CLIP | HRAimp | 91.38 | 95.18 | 91.34 | 95.34 | 97.21 | 98.17 | 78.52 | 91.93 | 85.81 | 93.16 | 68.06 | 83.33 | 75.48 | 88.79 |
| AdvCLIP | 93.87 | 95.07 | 42.46 | 45.53 | 16.12 | 23.44 | 26.82 | 37.27 | 23.04 | 27.41 | 15.26 | 16.36 | 29.60 | 29.52 | |
| SGA | 97.06 | 97.83 | 50.02 | 55.55 | 23.92 | 29.28 | 31.90 | 39.02 | 29.91 | 33.66 | 16.68 | 18.64 | 18.61 | 21.05 | |
| ETU | 96.64 | 97.3 | 61.26 | 64.95 | 25.79 | 31.70 | 31.63 | 40.27 | 30.52 | 33.09 | 17.99 | 20.00 | 20.97 | 23.10 | |
| FD-UAP | 83.12 | 89.63 | 26.40 | 32.17 | 17.09 | 22.68 | 28.46 | 36.74 | 21.52 | 27.47 | 13.20 | 16.43 | 15.82 | 19.68 | |
| C-PGC | 95.79 | 96.99 | 87.13 | 90.12 | 62.80 | 72.42 | 65.62 | 77.77 | 57.65 | 67.67 | 42.36 | 54.75 | 47.78 | 59.18 | |
| HRAran | 97.38 | 98.11 | 82.96 | 88.28 | 48.72 | 62.85 | 51.35 | 65.31 | 48.95 | 64.21 | 41.35 | 51.95 | 41.59 | 51.53 | |
| CLIP | HRAimp | 98.77 | 98.74 | 91.79 | 95.22 | 72.64 | 81.72 | 69.17 | 81.64 | 70.74 | 82.21 | 57.57 | 71.76 | 61.80 | 73.29 |
| AdvCLIP | 39.76 | 45.36 | 27.09 | 32.51 | 22.28 | 26.42 | 32.20 | 39.77 | 25.83 | 30.04 | 68.68 | 64.60 | 56.27 | 50.19 | |
| SGA | 36.90 | 41.82 | 23.54 | 30.94 | 18.58 | 24.37 | 30.87 | 38.20 | 22.36 | 28.20 | 85.51 | 82.43 | 32.54 | 27.62 | |
| ETU | 49.90 | 56.11 | 40.42 | 45.16 | 25.6 | 31.95 | 33.57 | 41.47 | 34.38 | 37.32 | 82.99 | 85.19 | 50.90 | 42.04 | |
| FD-UAP | 39.64 | 45.24 | 28.16 | 34.59 | 19.57 | 26.15 | 31.13 | 39.24 | 26.71 | 31.48 | 69.32 | 68.74 | 39.02 | 33.55 | |
| C-PGC | 68.78 | 74.95 | 64.20 | 71.83 | 57.19 | 65.07 | 62.95 | 71.76 | 58.49 | 66.09 | 91.88 | 87.19 | 77.33 | 71.94 | |
| HRAran | 73.60 | 80.76 | 68.61 | 67.46 | 55.74 | 64.91 | 56.96 | 69.03 | 60.47 | 67.26 | 90.85 | 91.02 | 70.34 | 65.71 | |
| ALBEF | HRAimp | 82.26 | 88.29 | 78.38 | 84.67 | 69.52 | 77.54 | 70.74 | 80.86 | 71.16 | 78.77 | 94.15 | 94.50 | 83.20 | 84.63 |
| AdvCLIP | 47.32 | 54.57 | 36.04 | 42.29 | 24.80 | 29.74 | 33.31 | 41.69 | 29.11 | 32.26 | 45.13 | 37.25 | 93.89 | 89.00 | |
| SGA | 37.39 | 45.86 | 28.36 | 33.89 | 21.14 | 25.57 | 31.13 | 39.46 | 26.44 | 30.97 | 30.69 | 40.86 | 92.28 | 88.57 | |
| ETU | 53.25 | 57.99 | 42.30 | 46.58 | 26.40 | 31.90 | 33.88 | 41.78 | 31.48 | 33.83 | 45.42 | 37.93 | 93.28 | 89.45 | |
| FD-UAP | 38.58 | 44.07 | 25.58 | 31.58 | 20.49 | 26.06 | 29.65 | 38.01 | 27.24 | 30.42 | 31.92 | 28.46 | 83.97 | 78.65 | |
| C-PGC | 72.70 | 79.97 | 66.20 | 75.18 | 63.49 | 68.28 | 64.33 | 72.08 | 64.36 | 70.09 | 64.78 | 61.66 | 93.57 | 87.43 | |
| HRAran | 78.59 | 87.96 | 73.89 | 84.74 | 60.47 | 83.25 | 58.18 | 78.27 | 58.60 | 82.43 | 74.32 | 73.10 | 94.21 | 93.72 | |
| TCL | HRAimp | 86.80 | 94.17 | 83.16 | 93.43 | 72.80 | 93.23 | 71.42 | 90.63 | 71.16 | 92.59 | 78.99 | 88.09 | 96.32 | 96.22 |
@@ IV-A Settings
Implementation details. For the fundamental experiments, we set the perturbation magnitudes to for images and for texts. We further evaluate the proposed method under varying perturbation budgets. PGD is utilized to solve the optimization problem for the image attack, with the number of iterations and the step size as . The batch size is set as 16. . For , we employ the KL-divergence to quantify the discrepancy between two samples.
Baselines. We compare our approach with several state-of-the-art methods, including:
AdvCLIP [55], which generates UAPs by breaking neighbourhood relationships between images and texts;
SGA [20], which divides large-batch learning into several small-batch learning and aggregates gradients of small-batch data to enhance the gradient stability and avoid dropping into local optima;
ETU [47], which utilizes ScMix augmentation to increase input diversity and enhances the local utility of UAPs;
FD-UAP [32], which is proposed to destroy significant features while strengthening the less significant features at low-level layers;
C-PGC [11] uses conditional generators to learn adversarial image and text embeddings by enlarging their discrepancy from the original data while aligning them with the farthest samples. Text perturbations are obtained by searching for words in a corpus that are semantically similar to the learned adversarial text embeddings.
AdvCLIP, SGA, ETU, and FD-UAP focus solely on image-modality attacks, whereas C-PGC and the proposed HRA support attacks on both the image and text modalities.
| Source Model | CLIP | ALBEF | ||||||||
| B@4 | METEOR | ROUGE | CIDEr | SPICE | B@4 | METEOR | ROUGE | CIDEr | SPICE | |
IV-B Cross-model Transferability
@@
| Test Dataset | MSCOCO | ||||||||||||||
| AdvCLIP | 23.13 | 34.36 | 0.41 | 0.95 | 0.19 | 0.76 | 24.30 | 32.06 | 18.50 | 24.89 | 7.04 | 10.92 | 8.02 | 11.83 | |
| SGA | 34.57 | 42.08 | 25.87 | 32.28 | 19.08 | 23.85 | 14.97 | 26.06 | 22.97 | 28.21 | 16.94 | 22.23 | 12.14 | 13.02 | |
| ETU | 39.76 | 47.38 | 32.04 | 38.04 | 27.70 | 32.41 | 35.75 | 43.31 | 30.68 | 33.85 | 34.49 | 49.26 | 27.57 | 24.64 | |
| FD-UAP | 36.62 | 42.38 | 22.68 | 30.79 | 17.36 | 21.38 | 30.10 | 37.39 | 21.06 | 26.61 | 41.49 | 48.88 | 22.80 | 21.88 | |
| C-PGC | 43.55 | 61.89 | 36.14 | 57.74 | 30.31 | 51.00 | 36.07 | 59.44 | 24.42 | 49.10 | 35.94 | 53.40 | 28.49 | 38.47 | |
| HRAran | 60.24 | 69.33 | 51.53 | 65.46 | 43.95 | 55.47 | 44.49 | 59.87 | 44.03 | 56.40 | 37.64 | 55.52 | 31.09 | 42.71 | |
| ALBEF | HRAimp | 73.93 | 81.58 | 66.29 | 79.25 | 60.05 | 71.66 | 62.34 | 74.97 | 61.58 | 71.30 | 46.51 | 66.46 | 41.31 | 57.71 |
Table II summarizes the attack performance on both image-to-text and text-to-image retrieval tasks, using various source models and the Flickr30K dataset for training. From these results, we draw the following observations.
@@ IV-B1 Comparison with baselines
All methods exhibit substantial performance degradation when transferred to unseen models and datasets, confirming that different models rely on distinct feature representations and thus present a significant model gap. Notably, UAPs remain highly effective on the same model even when evaluated on different datasets. For example, UAPs generated on CLIP transfer well from Flickr30K to MSCOCO. This indicates that existing attacks tend to produce perturbations that are strongly model-specific.
The inferiority of compared methods largely comes from insufficient consideration of characteristics of UAPs, including model- and data-agnostic natures, as well as cross-modal interactions. From the perspective of the image modality, AdvCLIP generates UAPs by destroying neighbourhood relationships in the source models, but fails to account for this in the target model. FD-UAP focuses on capturing model-agnostic cues from low-level features, but fails to exploit cross-modal interactions, limiting its transferability. In contrast, our disentangled adversarial example refinement strategy, i.e., incorporating ScMix augmentation and local utility enhancement, effectively explores cross-modal alignment and captures the data-independent nature of UAPs. It enables more reliable disruption of cross-modal relationships across models (as demonstrated by the results of our conference version, ETU, and the proposed HRA).
In addition, C-PGC produces both image and textual UAPs conditioned on the training distribution. This makes it rely on model-dependent features, limitsing its ability to transfer to unseen models. SGA attempts to alleviate overfitting by aggregating gradients from other mini-batches. However, aggregating gradients from many mini-batches may dilute the relevance of the update direction, which explains its limited improvements, particularly on CLIP and ALBEF in Table II. In comparison, our method utilizes both previous and estimated future gradients to regularize the current gradient, providing more informative update guidance and achieving superior transferability.

### v2

@@ III-A Preliminaries
The proposed HRA learns universal adversarial perturbations (UAPs) for both image and text modalities in a black-box setting, where target VLP models, datasets, and downstream tasks are inaccessible during training. To this end, we adopt a transfer-based attack framework that leverages a surrogate model and dataset to generate UAPs transferable across diverse models, datasets, and tasks.
Our approach is generic and imposes no assumptions on the surrogate model. For example, some VLP models employ a multimodal encoder [13, 35], while others do not [26]. To ensure broad generality, we uniformly utilize the unimodal encoders available in each VLP model, i.e., the image encoder and the text encoder, denoted as and , respectively. Let denote a multimodal training dataset, where is an image-text pair and is the number of pairs. The objective is to learn UAPs that can mislead other VLP models to incorrectly associate images and texts in a target dataset at the reference time. For the image modality, we learn UAPs constrained by an norm to ensure imperceptibility, i.e., . Here, represents the perturbation magnitude. For the text modality, due to token discreteness, we implement a universal attack by consistently substituting a word in the text with a replacement word, with imperceptibility controlled by a substitution budget .
III-B Overview
To enhance adversarial transferability, we propose Hierarchical Refinement Attack (HRA), a universal multimodal adversarial framework. For the image modality, HRA directly regularizes the optimization trajectory, as overfitting arises when updates converge to local optima. By incorporating both past and predicted future gradients, it stabilizes update directions and prevents premature convergence. For the text modality, HRA hierarchically ranks word importance by masking individual tokens and measuring the divergence between masked and original representations, accounting for both intra- and inter-sentence contexts. An illustration of HRA is shown in Fig. 1.
III-C Future-aware Momentum
VLP models commonly bridge the gap between heterogeneous data, i.e., images and texts, by creating a shared representation space in which semantically related visual and linguistic inputs are closely aligned. This makes it natural for an attacker to disturb the alignment between paired images and texts. Specifically, given a surrogate model, A typical method for learning image UAPs is to enlarge the feature distance between matched images and texts:
@@ III-C Future-aware Momentum
Iterative gradient-based optimization methods such as PGD [11, 22] are commonly used to solve the above objective. However, under multi-round optimization, the update process tends to converge to suboptimal local minima, leading to overfitting [5, 32, 30, 15]. Momentum-based methods have been shown to stabilize update directions by accumulating historical gradients [20, 10]. Classical momentum regularizes the optimization trajectory solely based on past gradients. In the context of universal adversarial perturbation learning, where gradients are aggregated across the entire dataset, historical gradients may become misaligned with the current optimization direction, limiting attack effectiveness. Moreover, conventional momentum operates within a restricted temporal scope, resulting in a narrow exploration space. To address these limitations, we propose hierarchical future-aware momentum, which additionally incorporates estimated future gradients to regularize the optimization trajectory. This expanded temporal hierarchy stabilizes updates, broadens the search space, and effectively mitigates overfitting.
@@
We regularize the current gradient using both the previous and future gradients as follows:
@@ III-D Text Attack
The detailed algorithm is summarized in Algorithm 1.
@@ IV-A Settings
Implementation details. For the fundamental experiments, we set the perturbation magnitudes to for images and for texts. We further evaluate the proposed method under varying perturbation budgets. PGD is utilized to solve the optimization problem for the image attack, with the number of iterations and the step size as . The batch size is set as 16. . For , we employ the KL-divergence to quantify the discrepancy between two samples. We follow the data augmentation strategy introduced in [42].
@@
| Source Model | CLIP | ALBEF | ||||||||
| B@4 | METEOR | ROUGE | CIDEr | SPICE | B@4 | METEOR | ROUGE | CIDEr | SPICE | |
Baselines. We compare our approach with several state-of-the-art methods, including AdvCLIP [50], SGA [17], ETU [42], FD-UAP [29], C-PGC [8]. AdvCLIP, SGA, ETU, and FD-UAP focus solely on image-modality attacks, whereas C-PGC support attacks on both the image and text modalities.
IV-B Cross-model Transferability
@@
Table I summarizes the attack performance on both image-to-text and text-to-image retrieval tasks, using various source models and the Flickr30K dataset for training. From these results, we draw the following observations.
@@ IV-B1 Comparison with baselines
All methods exhibit substantial performance degradation when transferred to unseen models and datasets, confirming that different models rely on distinct feature representations and thus incur a significant model gap. From the image-modality perspective, AdvCLIP generates UAPs by destroying neighbourhood relationships in the source models, but fails to account for this in the target model. FD-UAP focuses on low-level features, assuming them to be more transferable. However, this would also lead to overfitting to low-level features specific to the source model. C-PGC produces both image and textual UAPs conditioned on the training distribution. This makes it rely on model-dependent features, limiting its ability to transfer to unseen models. SGA attempts to alleviate overfitting by aggregating gradients from other mini-batches. However, aggregating gradients from many mini-batches may dilute the relevance of the update direction, which explains its limited improvements, particularly on CLIP and ALBEF in Table I. In contrast, our method utilizes both historical and estimated future gradients to regularize the current gradient, providing more informative optimization guidance and reducing the risk of convergence to local optima. Consequently, it achieves superior transferability across models.